# RATE: Causal Explainability of Reward Models with Imperfect Counterfactuals

### v2

@@ Abstract
Reward models are widely used as proxies for human preferences when aligning or evaluating LLMs. However, reward models are black boxes, and it is often unclear what, exactly, they are actually rewarding. In this paper we develop Rewrite-based Attribute Treatment Estimator (RATE) as an effective method for measuring the sensitivity of a reward model to high-level attributes of responses, such as sentiment, helpfulness, or complexity. Importantly, RATE measures the causal effect of an attribute on the reward. RATE uses LLMs to rewrite responses to produce imperfect counterfactual examples that can be used to measure causal effects. A key challenge is that these rewrites are imperfect in a manner that can induce substantial bias in the estimated sensitivity of the reward model to the attribute. The core idea of RATE is to adjust for this imperfect-rewrite effect by rewriting twice. We establish the validity of the RATE procedure and show empirically that it is an effective estimator. Code is available at https://github.com/toddnief/RATE.
@@ 1 Introduction
Reward models (RMs) play a critical role in aligning large language models (LLMs) with desired behaviors. These models evaluate the quality of LLM outputs and are widely used in, e.g., post-training, inference-time sampling adjustment (through best-of-$n$ sampling), and LLM evaluation. However, reward models, typically implemented as fine-tuned LLMs, are black boxes. Accordingly, it can be difficult to understand what a reward model is actually rewarding. Despite their importance, explainability research for RMs remains relatively underexplored (Lambert et al., 2024).
To understand reward models we would like to be able to measure their responsiveness to high-level attributes of responses, such as helpfulness, correctness, or sentiment. Such responsiveness measurements would serve as an important diagnostic tool. The purpose of the present paper is to develop such a measurement procedure.
Naively, one might attempt to estimate RM responsiveness to an attribute by comparing the average reward assigned to responses with and without the attribute using a labeled dataset. However, this approach is flawed: it conflates the target attribute’s influence with correlations present in the evaluation data. For instance, suppose we are measuring how much the RM responds to “sentiment” and, unbeknownst to us, highly-negative samples in our evaluation data tend to have more typos and sloppy formatting. Then if we naively average the reward over the negative samples, and again over the positive samples, we’ll inadvertently measure the effect of not just sentiment, but typos and formatting as well. In particular, this means we can measure a large effect of an attribute on the RM, even if the RM is completely insensitive to that attribute.
@@
There is a growing literature on estimating the causal effects of attributes of text (Feder et al., 2022; Grimmer et al., 2022; Jin et al., 2022; Chen & Chu, 2023; Gui & Veitch, 2023). Generally, this work provide methods for estimating the causal effect by either making strong assumptions on the evaluation data or compensating for potential confounding factors.
Notice that if we had access to counterfactual pairs of responses (i.e., pairs where the only difference is the attribute of interest), we could estimate the target effect by simply comparing the rewards assigned to each response. A natural idea is to use LLMs to generate such pairs by rewriting responses to change the target attribute. If the rewrites were perfect, we could directly measure the causal effect of the attribute on the reward. However, in practice, LLMs produce imperfect rewrites, changing off-target attributes as well. These imperfections can substantially bias the estimated causal effect, as we will see in Section 5.
@@
    We empirically demonstrate that RATE correctly estimates causal effects, while the naive and single-rewrite baselines are highly influenced by spurious associations. We also showt that this causal-vs-correlational problem substantially affects the explanations of leading reward models using common benchmarks.
@@ 2 Setup
Figure 1: Correlations make it hard to isolate the effect of individual attributes on reward models. For instance, if we wish to measure how much a reward model responds to ‘helpfulness’, we must isolate this from correlated attributes (e.g. length).
@@
If we want to measure the sensitivity of a given reward model to an attribute of interest such as helpfulness, the obvious approach is to take the dataset of prompt-completion pairs, label each completion as helpful or unhelpful, then check whether the rewards for the helpful responses are higher than the rewards for the unhelpful responses. Mathematically, we define this average conditional reward difference as:
@@
where the expectation is taken over the distribution from which our evaluation examples are drawn. The problem here is that, even in the infinite data limit, this quantity does not generally isolate the effect of $W$ on $R$. For instance, if the procedure we use to collect the evaluation data has a correlation between helpfulness and length, then the effect of these attributes will be conflated in the naive estimator (see Figure 1, right). Consequently, we will misinterpret the true behavior of the reward model.
@@
Intuitively, if $W=1$ is a helpful response, the ATT measures the change in reward when we take a helpful response and make it less helpful, and the ATU measures the change in reward when we take an unhelpful response and make it more helpful. These estimands can differ substantially from each other and from the ATE (see Figure 8 in Section D.1). There is no reason to expect these quantities to align in general, so thought should be given to which is most relevant to the question at hand. Indeed, even human preferences are often asymmetric (Kahneman & Tversky, 2013), so we might expect reward model preferences to be as well.
@@
Table 1: Recent language model show promise as (potentially imperfect) rewriters, since qualitatively they are capable of following instructions to change the target attribute (even if they do not always leave off-target attributes unchanged). For instance, here GPT-4o qualitatively does well at rewriting IMDB responses to change sentiment from negative (W = 0) to positive (W = 1). However, this does not ensure that GPT-4o will not change other attributes beside sentiment.
@@ 3 Rewrite-based Attribute Treatment Estimator (RATE)
We use $\text{Re}(x^{i},y^{i},w)$ to denote the rewrite operation, which takes a prompt-response pair $(x^{i},y^{i})$ and a desired attribute value $w$, returning a modified response $\tilde{y}^{i}$ such that $W(x^{i},\tilde{y}^{i})=w$. When $x^{i}$ is clear from context, we drop it for brevity as $\text{Re}(y^{i},w)$, even though the rewriter in general may depend on the input prompt.
@@
This is inevitable. Ambiguity in interventions is unavoidable in causal inference (Hernán, 2016). In our context, there is subjectivity in what helpfulness, complexity, or sentiment actually mean. An advantage of rewrite instructions is that we can use natural language to specify, as clearly as possible, what property we are trying to modify. We can understand whether our instructions are having the intended effect by qualitatively examining the rewritten outputs and checking that they indeed change the target attribute. In practice, finding effective rewrite instructions requires an iterative cycle of generating rewrites, examining the responses, and adjusting the rewrite prompt to be more clear and specific.
@@
For example, in Table 1 (and Section E.3), sentiment flips successfully, but other properties may also shift. This issue can be identified quantitatively by comparing reward distributions before and after rewriting (Figure 2). Ideally, if only sentiment changed, a second rewrite reversing the attribute should restore the original distribution. Otherwise, the (original, rewrite) pairs are not perfectly counterfactual and cannot directly estimate Equation 1.
Mathematically, rewriting introduces some error $\epsilon^{i}_{w}$ in the observed reward because we cannot perfectly construct the true counterfactual $y^{i}(w)$, which should only differ in the target attribute:
$\epsilon^{i}_{w}=R(x^{i},\text{Re}(y^{i},w))-R(x^{i},y^{i}(w))$
@@ 4 Cancellation of Multiple Rewrite Errors
We now turn to the validity of the RATE procedure. Intuitively, the idea is that as long as the errors introduced by the rewrite are independent of the target attribute $W$, then these errors will cancel out when we compute the average. A particularly important special case of independence is when the LLM always introduces the same error, irrespective the rewrite direction. For example, we observe that GPT-4o rewrites nearly always correct typos, independent of the attribute being rewritten. In this case, the single-rewrite estimator is biased because it includes the effect of the typo correction, but RATE is not because it compares typo-corrected responses to typo-corrected responses. We now formalize these intuitions.
@@
Continuing the example where rewriting sentiment ($W$) corrects typos ($\xi$) but leaves language ($Z$) unchanged, this states that GPT-4o isn’t more or less likely to fix typos on positive sentiment responses compared to negative sentiment. Note that this is in fact weaker than assuming that LLM-rewriters tend to repeat the same rewrite errors (e.g. GPT-4o always fixes typos no matter what). If we do observe this behavior, then the assumption is trivially satisfied.
@@
Intuitively, this assumption is saying that at the level of the reward model there aren’t any interactions between the component corresponding to potential rewrite-errors $\xi$ and either $W$ or $Z$.
@@
See Appendix B for the proof.
@@ 5 Experiments
Answering the first question requires knowing ground truth causal effects. To this end, we design semi-synthetic experiments with known ground truth. Using these, we find that RATE is effective at estimating the true effects, while the naive and single-rewrite estimators fail. The second question is whether the correctness of the RATE estimator over the naive estimator actually matters in practice. We find that across a variety of reward models, attributes, and datasets, the RATE estimates differ substantially from the naive baseline (see Figure 5). Accordingly, the distinction between RATE and the naive estimator is indeed important. As a particular example, the RATE estimator shows that the phenomenon of “length bias” in rewards may actually be mainly an artifact of bias in naive evaluations.
@@ 5.1 Semi-synthetic Experiments
As seen in Table 2, GPT-4o corrects typos even when not asked to, which we predict introduces bias for the single-rewrite estimator. By artificially correlating typos with “starts with a vowel” in the dataset, the single-rewrite procedure ought to demonstrate a positive bias to the change in reward score when examples are rewritten to start with a consonant (since typos harm the reward score). This results in a negatively biased estimate of the treatment effect of “starts with a vowel”—the estimation procedure views “starts with a vowel” as harming the reward score of an example.
@@ 7 Conclusion and Discussion
RATE only addresses the effect of the attribute on the reward model, without reference to the downstream task that the reward model will be used for. This is both a strength and a limitation. The strength is that it allows us to understand the reward model in isolation, which is significant since a given reward model may be used for many tasks. The limitation is that it is not completely clear how the causal sensitivity the reward model will translate to, e.g., the behavior of a LLM aligned to this reward. It would be an interesting direction for future work to elucidate this connection.

### v3

@@ Abstract
Reward models are widely used as proxies for human preferences when aligning or evaluating LLMs. However, reward models are black boxes, and it is often unclear what they are actually rewarding. In this paper, we develop Rewrite-based Attribute Treatment Estimator (RATE) as an effective method for measuring the sensitivity of a reward model to high-level attributes of responses, such as sentiment, helpfulness, or complexity. Importantly, RATE measures the causal effect of an attribute on the reward. RATE uses LLMs to rewrite responses to produce imperfect counterfactual examples that can be used to measure causal effects. A key challenge is that these rewrites are imperfect in a manner that can induce substantial bias in the estimated sensitivity of the reward model to the target attribute. The core idea of RATE is to adjust for this imperfect-rewrite effect by rewriting twice. We establish the validity of the RATE procedure and show empirically that it is an effective estimator. Code is available at https://github.com/toddnief/RATE.
@@ 1 Introduction
Reward models (RMs) play a critical role in aligning large language models (LLMs) with desired behaviors. These models evaluate the quality of LLM outputs and are widely used in, e.g., post-training, inference-time sampling adjustment (through best-of-$n$ sampling), and LLM evaluation. However, reward models, typically implemented as fine-tuned LLMs, are black boxes. Accordingly, it can be difficult to understand what a reward model is actually rewarding. Despite their importance, explainability research for RMs remains underexplored (Lambert et al., 2024).
To understand reward models, we would like to be able to measure their responsiveness to high-level attributes of responses, such as helpfulness, correctness, or sentiment. The ability to measure the responsiveness of an RM to an attribute would serve as an important diagnostic tool, and the purpose of this paper is to develop such a measurement procedure.
Naively, one might attempt to estimate RM responsiveness to an attribute by comparing the average reward assigned to responses with and without the attribute, using a labeled dataset. However, this approach is flawed: it conflates the target attribute’s influence with correlations present in the evaluation data. For instance, suppose we are measuring how much an RM responds to “sentiment” and, unbeknownst to us, highly-negative samples in our evaluation data tend to have more typos and sloppy formatting. Then if we naively average the reward over the negative samples, and again over the positive samples, we’ll inadvertently measure the effect of not just sentiment, but typos and formatting as well. In particular, this means we can measure a large effect of an attribute on the RM, even if the RM is completely insensitive to that attribute.
@@
Notice that, if we had access to counterfactual pairs of responses (i.e., pairs where the only difference is the attribute of interest), we could estimate the target effect by simply comparing the rewards assigned to each response. A natural idea is to use LLMs to generate such counterfactualpairs by rewriting responses to change only the target attribute. If the rewrites were perfect, we could directly measure the causal effect of the attribute on the reward. However, in practice, LLMs produce imperfect rewrites, changing off-target attributes as well. These imperfections can substantially bias the estimated causal effect, as we will see in Section 5.
Figure 1: When generating counterfactual pairs, LLMs change other attributes, such as tone, length, or grammar. Empirically, using the rewrites of rewrites corrects for this bias. (Left) Naively sampling pairs which differ on the attribute of interest (e.g., sentiment) will lead to a biased estimate of the causal effect because other attributes may also change. (Middle) When we rewrite a response to change the attribute of interest (e.g., from positive to negative sentiment), the LLM may also change other attributes, such as fixing typos. (Right) Rewriting the rewritten response again tends to cancel out these off-target changes, in a manner we make precise in Section 3.
There is a growing literature on estimating the causal effects of attributes of text (Feder et al., 2022; Grimmer et al., 2022; Jin et al., 2022; Chen & Chu, 2023; Gui & Veitch, 2023). Generally, this work provide methods for estimating causal effects by either making strong assumptions on the evaluation data or compensating for potential confounding factors.
Note that our question fundamentally differs from the similarly named counterfactual explanation paradigm (Wachter et al., 2018) studied recently in the context of reward models by Jiang et al. (2024). While counterfactual explanations aim to explain the behavior of a model by finding a minimally perturbed input that changes the model’s output, we are interested in measuring the average treatment effect of an attribute on the assigned reward, as measured across many counterfactual pairs.
@@
    We empirically demonstrate that RATE correctly estimates causal effects, while the naive and single-rewrite baselines are highly influenced by spurious associations. We also show that this causal-vs-correlational problem substantially affects the explanations of leading reward models using common benchmarks.
@@ 2 Setup
If we want to measure the sensitivity of a given reward model to an attribute of interest, such as helpfulness, the obvious approach is to take the dataset of prompt-completion pairs, label each completion as helpful or unhelpful, then check whether the rewards for the helpful responses are higher than the rewards for the unhelpful responses. Mathematically, we define this average conditional reward difference as:
@@
where the expectation is taken over the distribution from which our evaluation examples are drawn. The problem here is that, even in the infinite data limit, this quantity does not generally isolate the effect of $W$ on $R$. For instance, if the procedure we use to collect the evaluation data has a correlation between helpfulness and length, then the effect of these attributes will be conflated in the naive estimator (see Figure 1, middle). Consequently, we will misinterpret the true behavior of the reward model.
@@
Intuitively, if $W=1$ is a helpful response, the ATT measures the change in reward when we take a helpful response and make it unhelpful, and the ATU measures the change in reward when we take an unhelpful response and make it helpful. These estimands can differ substantially from each other and from the ATE (see Figure 8 in Section D.1). There is no reason to expect these quantities to align in general, so thought should be given to which is most relevant to the question at hand. Indeed, even human preferences are often asymmetric (Kahneman & Tversky, 2013), so we might expect reward model preferences to be as well.
@@
Table 1: Recent language models show promise as (imperfect) rewriters, since, qualitatively, they are capable of following instructions to change a target attribute (even if they do not always leave off-target attributes unchanged). For instance, GPT-4o qualitatively does well at rewriting IMDB responses to change sentiment from negative (W = 0) to positive (W = 1). However, this does not ensure that GPT-4o will not change other attributes besides sentiment.
@@ 3 Rewrite-based Attribute Treatment Estimator (RATE)
We use $\text{Re}(x^{i},y^{i},w)$ to denote the rewrite operation, which takes a prompt-response pair $(x^{i},y^{i})$ and a desired attribute value $w$, returning a modified response $\tilde{y}^{i}$ such that $W(x^{i},\tilde{y}^{i})=w$. When $x^{i}$ is clear from context, we drop it for brevity and write the operation as $\text{Re}(y^{i},w)$, even though the rewriter in general may depend on the input prompt.
@@
Ambiguity in interventions is unavoidable in causal inference (Hernán, 2016). In our context, there is subjectivity in what helpfulness, complexity, or sentiment actually mean. An advantage of rewrite instructions is that we can use natural language to specify, as clearly as possible, what property we are trying to modify. We can understand whether our instructions are having the intended effect by qualitatively examining the rewritten outputs and checking that they indeed change the target attribute. In practice, finding effective rewrite instructions requires an iterative cycle of generating rewrites, examining the responses, and adjusting the rewrite prompt to be more clear and specific.
@@
For example, in Table 1 (and Section E.3), sentiment flips successfully, but other properties may also shift. This issue can be identified quantitatively by comparing reward distributions before and after rewriting (Figure 2). Ideally, if only sentiment changed during rewriting, a second rewrite reversing the attribute should restore the original distribution. Otherwise, the (original, rewrite) pairs are not perfectly counterfactual and cannot directly estimate Equation 1.
Mathematically, rewriting introduces some error $\epsilon^{i}_{w}$ in the observed reward because we cannot perfectly construct the true counterfactual $y^{i}(w)$, which differs only on the target attribute:
$\epsilon^{i}_{w}=R(x^{i},\text{Re}(y^{i},w))-R(x^{i},y^{i}(w))$ (2)
@@ 4 Cancellation of Multiple Rewrite Errors
We now turn to the validity of the RATE procedure. Intuitively, the idea is that, as long as the errors introduced by the rewrite are independent of the target attribute $W$, then these errors will cancel out when we compute the average. A particularly important special case of independence is when the LLM always introduces the same error, irrespective the rewrite direction. For example, we observe that GPT-4o rewrites nearly always correct typos, independent of the attribute being rewritten. In this case, the single-rewrite estimator is biased because it includes the effect of the typo correction. RATE, however, is not, because it compares typo-corrected responses to typo-corrected responses. We now formalize these intuitions.
@@
Continuing the example where rewriting sentiment ($W$) corrects typos ($\xi$) but leaves language ($Z$) unchanged, this states that GPT-4o isn’t more or less likely to fix typos on positive sentiment responses compared to negative sentiment. Note that this is, in fact, weaker than assuming that LLM-rewriters tend to repeat the same rewrite errors (e.g. GPT-4o always fixes typos no matter what). If we do observe this behavior, then the assumption is trivially satisfied.
@@
Intuitively, this assumption is saying that, at the level of the reward model, there aren’t any interactions between the component corresponding to potential rewrite-errors $\xi$ and either $W$ or $Z$.
@@
See Appendix B for the proof. We emphasize that these are merely sufficient conditions to show that there exist any situations under which RATE is consistent. There’s no reason a priori to expect that imperfect rewrites can provide a causal estimation. Hence, the purpose of the theorem is simply to show that the approach is not vacuous.
@@ 5 Experiments
Answering the first question requires knowing ground truth causal effects. To this end, we design semi-synthetic experiments with known ground truth. In this setting, we find that RATE is effective at estimating the true effects, while the naive and single-rewrite estimators fail.
We note that a quantitative evaluation, as outlined in Section 5.1, can be performed to determine whether rewrite errors cancel out during estimation with the double-rewrite estimator. If they do not, sensitivity to correlations between on-target and off-target attributes would be observed.
The second question is whether the correctness of the RATE estimator over the naive estimator actually matters in practice. We find that across a variety of reward models, attributes, and datasets, the RATE estimates differ substantially from the naive baseline (see Figure 5). Accordingly, the distinction between RATE and the naive estimator is indeed important. As a particular example, the RATE estimator shows that the phenomenon of “length bias” in rewards may actually be mainly an artifact of bias in naive evaluations.
@@ 5.1 Semi-synthetic Experiments
As seen in Table 2, GPT-4o corrects typos even when not asked to, which, we predict, introduces bias for the single-rewrite estimator. By artificially correlating typos with “starts with a vowel” in the dataset, the single-rewrite procedure ought to demonstrate a positive bias to the change in reward score when examples are rewritten to start with a consonant (since typos harm the reward score). This results in a negatively biased estimate of the treatment effect of “starts with a vowel”—the estimation procedure views “starts with a vowel” as harming the reward score of an example.
@@ 7 Conclusion and Discussion
RATE only addresses the effect of the attribute on the reward model, without reference to the downstream task that the reward model will be used for. This is both a strength and a limitation. The strength is that it allows us to understand the reward model in isolation, which is significant since a given reward model may be used for many tasks. The limitation is that it is not completely clear how the causal sensitivity the reward model will translate to, e.g., the behavior of a LLM aligned to this reward. Exploring this relationship further could provide valuable insights into the role of reward models in the alignment process.